The reaction between potassium chlorate and red phosphorus takes place when you strike a match on a matchbox. If you were to react 53.7 g of potassium chlorate (KClO3) with excess red phosphorus, what mass of tetraphosphorus decaoxide (P4O10) could be produced?
1
Expert's answer
2016-10-25T14:57:11-0400
10KClO 3 + 3P 4 = 3P 4 O 10 + 10KCl

n = m/M

M (KClO 3 ) = 122.6 g/mol

M (P 4 O 10 ) = 283.9 g/mol

n (KClO 3 ) = 53.7 / 122.6 = 0.4 mol

n (P 4 O 10 ) = n (KClO 3 )/10 · 3 = 0.4/10 · 3 = 0.12 mol

m (P 4 O 10 ) =n · M = 0.12 · 283.9 = 34.1 g